# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v2

@@ VI-B Minimizing E2E Energy Consumption
VII Numerical Results
In this section, we present numerical results to evaluate the performance of the proposed joint blocklength and power allocation algorithms. We consider a total area of 500 $\rm m\times$ 500 $\rm m$, with the sensing location positioned at the center. The $N_{\rm tx}=16$ ISAC transmit APs are uniformly distributed in the area. We consider the number of sensing receive APs equal to $N_{\rm rx}=1$ or $N_{\rm rx}=2$ where the first AP is located at $(200,250)$ and the second one is located at $(300,250)$. The number of antenna elements per AP is set to $M=4$. The total number of URLLC UEs in the network is $N_{\rm ue}=8$. We set $P_{\rm tx}=100$ mW and the uplink pilot transmission power for each UE is $50$ mW.
The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [43, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4]. The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to $0.3$ to suppress the known parts of the target-free channels due to LOS and permanent obstacles [34]. The sensing channel gains are computed by the bi-static radar range equation [9]. The carrier frequency, the bandwidth, and the noise variance are set to $1.9$ GHz, $200$ KHz, and $-114$ dBm, respectively. The number of pilot symbols is $L_{p}=10$.
The spatial correlation matrices for the communication channels are generated by using the local scattering model in [35, Sec. 2.5.3]. The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0\,\rm dBsm$. For all the UEs, the packet size, maximum transmission delay, and DEP threshold are $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{\rm th}=10^{-5}$, respectively. In the proposed algorithms, $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and $c_{\rm max}=30$. The sensing SINR threshold is set to $\gamma_{s}=0$ dB, unless otherwise stated. The remaining parameters are detailed in Table II, where the values are consistent with those in [30], except for a scaling of the total processing capacity and the corresponding power of a DU by a factor of ten to accommodate all the processing in a single DU.
TABLE II: Simulation Parameters
$\Delta^{\rm tr},\sigma_{\rm cool}$ 4, 0.9 $P_{\rm fixed}$ 120 W
@@ VII Numerical Results
$\Delta_{\rm cloud}^{\rm proc}$ 740 W $C_{\rm max}$ 1800 GOPS
Figure 2: (a) Transmission energy consumption, (b) E2E energy consumption, and (c) breakdown of energy consumption over one transmission for different optimization algorithms, for clutter-aware target detector with $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
We compare the performance of two proposed algorithms: i) SeURLLC+ aiming to minimize the transmission energy consumption, and ii) E2E SeURLLC+ aiming to minimize the E2E energy consumption. Additionally, we include two baseline algorithms for URLLC systems without sensing integration: (i) URLLC-only, which minimizes transmission energy consumption, and (ii) E2E URLLC-only, which minimizes E2E energy consumption. We further compare results using clutter-aware and clutter-unaware MAPRT detectors, as defined in (17) and (18), respectively.
Fig. 2(a) illustrates the transmission energy consumption for four algorithms with the clutter-aware detector. The results show that integrating sensing into the system necessitates higher transmission energy. Notably, the URLLC-only and SeURLLC+ achieve lower transmission energy compared to their E2E counterparts, as their primary objective is to minimize transmission energy consumption.
The E2E energy consumption is depicted in Fig. 2(b). We observe that E2E energy minimization results in a substantial reduction in total energy consumption by approximately $50\%$ in URLLC-only scenarios, while its impact is less pronounced in ISAC scenarios, resulting in only a marginal decrease. The E2E algorithms try to minimize the total energy consumption by reducing the blocklength. However, higher power coefficients are required to meet the reliability requirements for URLLC UEs. In SeURLLC+, higher power coefficients for sensing allow lower blocklength, narrowing the gap in optimal blocklengths between SeURLLC+ and E2E SeURLLC+.
Fig. 2c provides a detailed breakdown of energy consumption by system components and different operations, including ISAC transmit APs, sensing receive APs, communication and sensing processing, and “Others,” which represent load-independent and idle mode power consumption in the cloud. For both algorithms, most of the energy consumption is attributed to sensing processing tasks, ISAC APs, and load-independent and idle mode power consumption in the cloud. However, the E2E SeURLLC+ algorithm slightly reduces the corresponding energy consumption.
Figure 3: (a) Blocklength, (b) E2E energy consumption, (c) detection probability, and (d) sensing update rate vs. decoding error probability threshold with clutter-aware target detector, $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
VII-A Impact of URLLC Requirements
Fig. 3 illustrates the blocklength, E2E energy consumption, detection probability and sensing update rate as functions of DEP threshold. Generally, the blocklength in Fig. 3(a) decreases as the reliability threshold is relaxed in all algorithms except for URLLC-only. The URLLC-only algorithm consistently uses the maximum available blocklength, i.e., $L_{\rm max}$. By integrating sensing, the optimal blocklength is significantly reduced compared to URLLC-only, as meeting sensing constraint requires higher power consumption resulting in lower blocklength. Relaxing the reliability constraint would generally reduce the E2E energy consumption. However, ISAC scenarios demand higher energy consumption as it is expected as shown in Fig. 3(b).
Fig. 3(c) and Fig. 3(d) demonstrate the impact of the DEP threshold on sensing performance. Relaxing the DEP constraint slightly reduces detection probability while increasing the sensing update rate, due to smaller blocklengths at higher DEP thresholds. Notably, the E2E SeURLLC+ algorithm outperforms the SeURLLC+ algorithm in both sensing and communication performance.
It is worth mentioning that higher delay threshold allows the system to operate with higher blocklength. However, it does not effect the results since the algorithm aims to minimize the energy consumption. Therefore, it still chooses a smaller blocklength.
VII-B Impact of Sensing Requirement
The impact of the sensing SINR threshold, $\gamma_{s}$, on the performance of the E2E SeURLLC+ algorithm is evaluated in Figs. 4a-c in terms of E2E energy consumption, detection probability, and sensing update rate, respectively. Comparing Fig. 4(a) with Fig. 4(b), we observe that both detection probability and energy consumption generally increase with a higher SINR threshold. The clutter-aware detector with $N_{\rm rx}=2$ provides the highest detection probability, consistently above 0.9, whereas the clutter-unaware detector cannot achieve a detection probability over 0.6, despite offering the lowest energy consumption. Notably, using a clutter-aware detector while disabling one RX-AP maintains a detection probability above 0.8 with around 43% energy saving. This suggests a potential optimization strategy: turning off one sensing receive AP can halve the energy consumption, significantly enhancing overall energy efficiency. Higher sensing SINR threshold demands higher blocklength, resulting in lower sensing update rate as shown in Fig. 4(c). Comparing Fig. 4(b) and Fig. 4(c), it is evident that increasing the sensing SINR threshold significantly decreases the sensing update rate without substantially improving detection probability.
Figure 4: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. sensing SINR threshold with E2E SeURLLC+ algorithm.
VII-C Impact of Number of APs
We explore the impact of the number of transmit and receive APs on the network performance in terms of E2E energy consumption, detection probability, and sensing update rate as shown in Figs. 5a-c, respectively. The results in Fig. 5(a) and Fig. 5(b) indicate that increasing the number of receive APs generally raises both E2E energy consumption and detection probability. For $N_{\rm rx}=2$, the detection probability substantially improves with only a marginal raise in energy consumption. However, further increasing to $N_{\rm rx}=4$ offers little improvement in detection probability while significantly increasing energy consumption. Moreover, the results in Fig. 5(c) show a consistent sensing update rate across all curves, especially for $N_{\rm tx}=8$ and $N_{\rm tx}=12$, indicating that $N_{\rm rx}$ has minimal impact on $R_{s}$. This suggests the optimization algorithm is more constrained by URLLC requirements than by sensing requirements.
A comparison between $N_{\rm tx}=8$ and $N_{\rm tx}=12$ in Fig. 5(c) reveals that a higher $N_{\rm tx}$ enhances URLLC satisfaction due to elevated SINR, resulting in a reduced blocklength and consequently a higher sensing update rate. It is observed that when $N_{\rm rx}=1$ or $N_{\rm rx}=2$, increasing the number of transmit APs does not significantly affect E2E energy consumption. This is because the rise in energy consumption from the additional APs is offset by the decrease in blocklength. Thus, the overall E2E energy consumption remains relatively unchanged. Conversely, when $N_{\rm rx}=4$, the influence of $N_{\rm tx}$ becomes more pronounced, as the energy consumption for sensing tasks increases non-linearly with the product of $N_{\rm tx}$ and $N_{\rm rx}$.
When $N_{\rm tx}=16$, the system’s behavior concerning E2E energy consumption and detection probability follows a similar trend, showing an increase in energy consumption and a decrease in detection probability. However, $R_{s}$ declines compared to when $N_{\rm tx}=12$, indicating a higher blocklength for $N_{\rm tx}=16$. This occurs because common power coefficients are used to maintain a favorable beam pattern, and increasing the number of transmit APs may constrain power due to some APs being closer to the UEs. As a result, a higher blocklength is necessary to fulfill the reliability requirements. These findings underscore the importance of optimizing the number of transmit and receive APs.
The results of detection probability is depicted in Fig. 5(b). In general, a decrease in blocklength translates to a lower detection probability explaining the results for $N_{\rm tx}=12$. However, we observe that detection probability for $N_{\rm tx}=16$ still decreases although blocklength is higher compared to $N_{\rm tx}=12$. This is because the power coefficients required to satisfy the sensing SINR requirement might be lower with higher $N_{\rm tx}$, resulting in a more pronounced negative impact on detection probability. This negative impact outweighs the benefits of increased blocklength, leading to a lower detection probability.
The findings emphasize the importance of optimizing the number of transmit and receive APs to balance energy consumption, detection probability and sensing update rate. Higher numbers of transmit APs can enhance URLLC satisfaction and reduce blocklength resulting in higher sensing update rate, but must be carefully managed to avoid an unnecessary increase in energy consumption and detection probability degradation.
Figure 5: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. the number of transmit ISAC APs for $N_{\rm rx}=1,2,4$ and E2E SeURLLC+ algorithm.
VIII Conclusion
In this paper, we proposed a joint blocklength and power control algorithm for a downlink CF-mMIMO system with multi-static sensing and URLLC UEs. We formulated two non-convex optimization problems focused on minimizing transmission and E2E energy consumption, and conducted GOPS analysis for both communication and sensing tasks. Numerical results demonstrated the effectiveness of the E2E energy minimization algorithm, especially in scenarios without sensing capabilities. Integrating sensing reduces the optimal blocklength compared to URLLC-only scenarios due to the higher power demands of sensing, which leads to a lower optimal blocklength. Moreover, the E2E algorithms minimize total energy consumption by further reducing the blocklength. Additionally, we observed that higher sensing SINR thresholds significantly lower the sensing update rate without substantially improving detection probability. Our findings also indicated that increasing the number of sensing receive APs raises sensing-related processing energy, suggesting that deactivating some APs can help reduce E2E energy consumption in CF-mMIMO ISAC systems. While increasing the number of transmit APs can enhance URLLC satisfaction and sensing update rate by reducing blocklength, careful management is necessary to prevent excessive energy consumption and degradation in detection probability.
@@ Appendix A Proof of Theorem 1
Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$. Then, minimizing the objective function is equivalent to minimizing the following convex function (quadratic-over-linear function) $\frac{\|\bm{\rho}\|^{2}}{\overline{L}}$ Minimizing this function, at the optimal solution, holds that $\overline{L}=1/L_{d}=1/(L-L_{p})$. The reliability constraints in (51b) can be written in the form of
@@
From (38), the sensing constraint in (51d) are expressed as
@@
Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (51e) in SOC form in terms of $\bm{\rho}$ as $\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ , for $k=1,\ldots,N_{\rm tx}$.

### v3

@@ VII-A Impact of Sensing and Communication Requirements
Fig. 5 illustrates the maximum blocklength threshold and network availability as a function of the refreshing rate threshold. The network availability represents the percentage of cases where the optimization problem is feasible, meaning that all the requirements can be satisfied [16]. As the required refreshing rate exceeds 1000 updates/s, the system becomes increasingly constrained, resulting in a notable reduction in allowable blocklength. Specifically, the maximum blocklength is nearly halved when $R_{\rm s}^{\rm th}=2000$. However, such reduced blocklengths may not satisfy reliability requirements, leading to infeasibility. Accordingly, the network availability drops to 80% at 1800 updates/s and further declines to 60% at 2000 updates/s.
Figs. 6a–c illustrate the impact of the DEP threshold on sensing performance, transmission energy, and total energy consumption, respectively. As shown in Fig. 6(a), stricter reliability requirements slightly enhance detection probability, since the communication task requires higher power and/or longer blocklengths to meet lower DEP thresholds, which in turn enhances the sensing performance. However, this improvement comes at the cost of increased energy consumption, as evident in Fig. 6(b) and Fig. 6(c). Notably, the E2E JPB-SeURLLC+ algorithm consistently outperforms the JPB-SeURLLC+ algorithm in terms of both sensing performance and energy consumption.
It is worth mentioning that higher delay threshold allows the system to operate with higher blocklength. However, it does not effect the results since the algorithm still chooses a smaller blocklength to minimize the energy consumption.
VII-B Impact of Detector and Number of APs
Fig. 7 presents a detailed breakdown of task-oriented energy consumption across system components and operations, including ISAC transmit APs, sensing receive APs, communication and sensing processing, and an “Others” category representing load-independent and idle-mode power consumption in the cloud. As shown, the majority of energy consumption is attributed to sensing processing tasks, ISAC transmit APs, and the cloud’s load-independent and idle-mode power usage.
To further explore the system behavior, Figs. 8a-c analyze the impact of blocklength, detector type, and the number of sensing receive APs on computational complexity, GOPS, and the number of required GPPs, respectively. This multi-dimensional analysis highlights how detector complexity and network configuration influence overall system efficiency. As shown in Fig. 8(a) and Fig. 8(b), clutter-aware detector with $N_{\rm rx}=2$ has the highest computational complexity and GOPS requirements among other configurations. Although longer blocklengths generally increase the number of operations for one task, they reduce the number of operations per unit time, thereby lowering GOPS requirements. This, in turn, reduces the number of active GPPs as shown in Fig. 8(c) and consequently decreases energy consumption through the $N_{\rm GPP}P_{\rm cloud,0}^{\rm proc}$ term in equation (32). Clutter-unaware detector maintains much lower complexity compared to the clutter-aware detector. However, a comparison between the clutter-aware and clutter-unaware curves reveals distinct trends. For the clutter-aware detector, the computational complexity remains approximately constant as blocklength increases, whereas for the clutter-unaware detector, complexity grows significantly with larger blocklengths. Interestingly, the opposite trend is observed in the GOPS curves. This behavior stems from the interplay between blocklength-dependent and blocklength-independent terms in the computational complexity expressions—specifically, equation (V-B) for the clutter-unaware detector and Table II for the clutter-aware detector. Based on these observations, we conclude that for the clutter-aware detector, blocklength-independent terms dominate the computational complexity, whereas for the clutter-unaware detector, blocklength-dependent terms are the primary contributors.
@@ VII-B Impact of Detector and Number of APs
VIII Conclusion and Future Directions
[…]
@@ Appendix A Proof of Theorem 1
Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$. Then, minimizing the objective function is equivalent to minimizing the convex function (quadratic-over-linear plus affine function) $\mathsf{F}$ in (44). Minimizing this function, at the optimal solution, leads to $\overline{L}=1/L_{d}=1/(L-L_{p})$. The reliability constraints in (40b) can be written in the form of
@@
From (27), the sensing constraint in (40e) are expressed as
@@
Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (40f) in SOC form in terms of $\bm{\rho}$ as $,\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ for $k=1,\ldots,N_{\rm tx}$.